to a 4.0e north of west. Vector B points 62.0 east of north. vector C points 11.0 west of south. These three of the following vectors vectors add to give a resuitant vector that is zero. Using components, find the Section 1.8 (a) due east (b) due north kom
Answer #1

A = 146 ( -cos34i + sin34 j) = - 121i + 81.6 j

B = b ( sin62 i + cos62 j)

C = c ( - sin11i - cos11 j)

A + B+ C = 0


(- 121 + b sin62 - c sin11)i + (81.6 + bcos62 - c cos11)j = 0


b sin62 - c sin11 = 121

bcos62 - c cos11 = - 81.6  


b = 173 units ...........Ans (A)


c = 165.8 units . .............Ans(B)


--------------------------

(A) due east = 292 cos30 =253 km


(B) due north = 292 sin30 = 146 km